# Benchmarking ECG Foundational Models: A Reality Check Across Clinical Tasks

### v1

@@ 4 Results
Table 2: Comparison of macro-AUROC (classification) and average z-normalized MAE (regression) under finetuning with linear prediction head. We use $\uparrow$$\downarrow$ to indicate whether higher or lower scores correspond to better model performance. Best results are bold and underlined; results not statistically worse are also bold. indicates label subsets only used for evaluation. Full results involving all considered models are given in Table 4. See Table 5 and Table 6 for frozen and linear evaluation. Model abbreviations: ECGFounder (Founder), ECG-JEPA (JEPA), ECG-FM (FM), ECG-CPC (CPC).
Foundation Models (Finetuned) Supervised
Founder JEPA ST-MEM MERL FM CPC S4
Adult ECG interpretation
PTB $\uparrow$ 0.656 0.679 0.694 0.717 0.725 0.702 0.654
Ningbo $\uparrow$ 0.974 0.973 0.954 0.955 0.971 0.973 0.972
CPSC2018 $\uparrow$ 0.966 0.974 0.946 0.936 0.972 0.969 0.962
CPSC-Extra $\uparrow$ 0.906 0.897 0.883 0.873 0.862 0.898 0.852
Georgia $\uparrow$ 0.920 0.918 0.888 0.912 0.912 0.913 0.903
Chapman $\uparrow$ 0.968 0.972 0.948 0.946 0.956 0.962 0.963
-Chapman (rhythm) $\uparrow$ 0.991 0.989 0.985 0.975 0.993 0.987 0.986
SPH $\uparrow$ 0.983 0.980 0.964 0.944 0.966 0.981 0.981
CODE-15% $\uparrow$ 0.987 0.991 0.974 0.982 0.986 0.989 0.991
PTB-XL (all) $\uparrow$ 0.934 0.940 0.908 0.925 0.927 0.949 0.941
-PTB-XL (diag) $\uparrow$ 0.950 0.946 0.904 0.942 0.926 0.951 0.943
-PTB-XL (form) $\uparrow$ 0.875 0.912 0.887 0.891 0.904 0.934 0.919
-PTB-XL (rhythm) $\uparrow$ 0.965 0.956 0.951 0.912 0.959 0.959 0.956
PTB-XL (sub) $\uparrow$ 0.943 0.935 0.916 0.937 0.932 0.940 0.938
PTB-XL (super) $\uparrow$ 0.935 0.921 0.901 0.930 0.916 0.934 0.932
Pediatric ECG interpretation
ZZU pECG $\uparrow$ 0.898 0.911 0.893 0.886 0.887 0.892 0.897
Cardiac structure & function
EchoNext (Echo) $\uparrow$ 0.817 0.817 0.816 0.822 0.772 0.831 0.819
Cardiac outcomes
MIMIC (Cardiac) $\uparrow$ 0.768 0.772 0.760 0.776 0.775 0.781 0.780
Non-cardiac outcomes
MIMIC (Non-cardiac) $\uparrow$ 0.701 0.711 0.688 0.712 0.719 0.719 0.714
Acute care predictions
MIMIC (Deterioration) $\uparrow$ 0.717 0.747 0.714 0.743 0.767 0.764 0.756
MIMIC (Mortality) $\uparrow$ 0.810 0.792 0.784 0.800 0.811 0.803 0.793
MIMIC (ICU) $\uparrow$ 0.748 0.742 0.737 0.744 0.750 0.753 0.745
Patient characteristics
MIMIC (Sex) $\uparrow$ 0.913 0.904 0.883 0.916 0.919 0.933 0.919
MIMIC (Age) $\downarrow$ 0.461 0.463 0.504 0.449 0.412 0.437 0.455
MIMIC (Biometrics) $\downarrow$ 0.637 0.640 0.673 0.625 0.620 0.604 0.626
MIMIC (ECG Features) $\downarrow$ 0.458 0.460 0.500 0.463 0.465 0.451 0.452
MIMIC (Lab Values) $\downarrow$ 0.679 0.677 0.688 0.676 0.688 0.673 0.675
MIMIC (Vital Signs) $\downarrow$ 0.704 0.703 0.715 0.702 0.704 0.700 0.701
@@
Table 2 shows an excerpt of finetuning results across tasks. Full finetuning, frozen evaluation, and linear evaluation results are given in Tables 4, 5, and 6 (Supplementary Material). Tables indicate statistically significant differences compared to the respective best method in this task. Ranked lists for all tasks and models are provided in Table 7. We further summarize this by reporting median ranks across all tasks of a given category in Figure 3. Appendix A.4 presents comparative predictions for the 10 best-performing conditions per task, illustrating both prediction quality and label diversity.
@@ 4.1 Supervised baselines
Supervised baseline vs. literature Our supervised baseline performs on par with or surpasses the literature results. The S4-based model achieves AUROCs of 0.941 (vs. 0.9417 (Mehari & Strodthoff, 2023)) on PTB-XL, 0.908 (vs. 0.843 (Strodthoff et al., 2024)) for cardiac discharge, 0.849 (vs. 0.764 (Strodthoff et al., 2024)) for non-cardiac discharge, 0.863 (vs. 0.752 (Alcaraz et al., 2025)) for clinical deterioration, 0.747 (vs. 0.746 (Alcaraz et al., 2025)) for ICU admission, and 0.874 (vs. 0.816 (Alcaraz et al., 2025)) for mortality. Improvements on MIMIC-based tasks (except for ICU) are attributed to the multi-task training objective.
@@
Impact of model architecture We also compare the supervised S4 model to the convolutional Net1D backbone used in ECGFounder, omitting supervised transformers, due to poor performance. Table 4 summarizes numerical results, and Table 7 shows statistically significant rankings: S4 consistently ranks first, while Net1D typically ranks at least four places lower. Although the backbone comparison was not the main goal, these results support the literature findings that CNNs are suboptimal for physiological time series (Strodthoff et al., 2024).
4.1.1 Finetuning
@@
Adult ECG interpretation: Across 11 tasks on 9 datasets, top performers are ECGFounder, ECG-JEPA, and ECG-CPC, often statistically surpassing the S4 baseline. ECG-FM ranks fourth overall, sometimes matching the top three but underperforming on Georgia, Chapman, and PTB-XL. MERL, ST-MEM, HuBERT-ECG, and ECGFM-KED generally fail to outperform the supervised baseline. Pediatric ECG interpretation: ECG-JEPA leads, followed by ECGFounder, ST-MEM, MERL, ECG-CPC, and S4, despite no pediatric pretraining data. Cardiac structure & function: ECG-CPC ranks first for echocardiography predictions, followed by ECGFounder, ECG-JEPA, ST-MEM, MERL, and S4. Cardiac and non-cardiac outcomes: ECG-CPC dominates, matching ECG-FM on non-cardiac and S4 on cardiac conditions. ECGFounder performs relatively poorly, likely due to little overlap between diagnostic labels during pretraining and target labels for the task. Acute care predictions: ECG-CPC and ECG-FM perform best across three tasks, followed by ECGFounder, ECG-JEPA, and MERL; none of them outperform S4 significantly. Patient characteristics: ECG-CPC ranks first in 5 of 6 tasks, outperforming S4 in 3. MERL and ECG-FM generally match or slightly underperform S4, while ECGFounder and ECG-JEPA typically fall below the supervised baseline.
4.2 Label efficiency
Figure 2: Scaling with dataset size on EchoNext (6 most-populated labels).
Label efficiency Scaling curves for three FMs are shown in Figure 2 with scaling parameters listed in Table 31. All models stay below the supervised baseline (S4) throughout the entire range of considered training set sizes. We use the parametric form of the fits to work out a label efficiency ratio $r=N^{*}/N$, i.e., the fraction of samples $N^{*}$ required for the pretrained model to reach the same performance as the supervised baseline $forgivenN$. For $N$ in the range of 250 to 1000, this yields label efficiency ratios between 0.30-0.62 for ECGFounder, 0.11-0.42 for ECG-JEPA and 0.21-0.40 for ECG-CPC, see Table 32 for details. This positions ECG-JEPA as the most label-efficient model, in particular in the very low sample size regime, closely followed by ECG-CPC. This establishes the label efficiency as a relevant benchmark parameter for foundation models. Furthermore, the results show that ECG FMs fulfill the promise of an improved label efficiency not only in comparison to the respective model architecture trained from scratch but also against a strong supervised baseline, improving label efficiency by up to a factor of 9.
@@ 4.3 Frozen and linear evaluation
Figure 3: Radar plots summarizing model performance ranks (lower rank indicates statistically significantly better performance) for the eight FMs and one of the supervised baselines (S4). The plot is based on data from Table 8.
@@ 5 Discussion
Details matter Reliable downstream performance depends on careful tuning. Layer-dependent learning rates consistently improve results, and some models (e.g., HuBERT-ECG) fail to train at all without them. Adjustable input size (where possible) also matters: using 2.5s crops with test-time averaging outperforms the full 10s inputs.
@@
Comparable assessment ECGFounder, ECG-JEPA, and ECG-CPC are strong on adult ECG interpretation, sometimes surpassing the supervised baseline. Pediatric ECG is dominated by ECG-JEPA. Other categories are led by ECG-CPC, followed by ECG-FM, while ECGFounder and ECG-JEPA are weaker here. Several models fall clearly below the supervised baseline, showing that self-supervised pretraining does not necessarily yield effective downstream performance. No single model consistently excels across all tasks and evaluation modes, though ECG-CPC comes closest to this goal with finetuning. This underscores that foundation model selection for downstream tasks should be informed by benchmarking results and similarity of the downstream task to the task categories considered here.
@@
Pretraining strategies This benchmark cannot definitively identify the optimal pretraining method, as downstream performance depends on the interplay of the pretraining methodology, dataset, and model architecture (Table 1). On the one hand, superior supervised performance may indicate more suitable models, which would favor SSM models over CNNs. On the other hand, transformers are naturally compatible with masking strategies that are commonly used in self-supervised methods. Larger datasets help in some cases (ECGFounder, ECG-CPC) but are not sufficient (e.g., HuBERT-ECG). Large-scale models do not automatically generalize well. ECG-CPC’s dominance outside ECG interpretation suggests self-supervised pretraining can be advantageous, though, its unidirectional backbone limits supervised performance compared to bidirectional models. These findings highlight the need for like-for-like comparisons of pretraining methods on a common dataset, such as HEEDB, using standardized architectures.
Model complexity Foundation models vary widely in complexity. This aspect is particularly relevant for clinical deployment or multimodal integration. Using the parameter count as an imperfect proxy (Dehghani et al., 2021), sizes range from 3.8M (ECG-CPC) to 97.2M (HuBERT-ECG). Notably, ECG-CPC matches ECGFounder and ECG-JEPA on adult ECG tasks with only 11% and 4% of their parameters respectively, and outperforms other FMs in most remaining categories.
@@
Model insights Frozen and linear evaluation reveal the nature of learned representations and the implicit knowledge captured by different foundation models. Insights from probing, as used here, remain coarse, but methods from explainable AI such as analyzing representation structures and their alignment across layers and models (Vielhaben et al., 2025) could provide more detailed insights into the knowledge acquired by these models.
@@
Limitations This work is subject to a number of limitations. First, the proposed tasks only includes in-distribution tests, out-of-distribution tests would require further datasets with comparable label sets. While this is often challenging for diagnostic labels due to mismatching ontologies, it is feasible for demographic data. Second, multi-tasking models seem to improve performance for classification tasks. In the regression case, we tried to alleviate this issue through z-normalization of the targets. Still the multi-tasking model used for computational efficiency might underperform for certain task compared to single-task models. Third, the consistent positive impact of layer-dependent learning rates during finetuning suggests the potential for developing improved finetuning methodologies.
@@ 6 Conclusion
In this work, we present a comprehensive benchmark for ECG foundation models across seven different task categories. Results showed heterogeneous performance across domains: three FMs (ECGFounder, ECG-JEPA and ECG-CPC) were found to exhibit strong performance in the most widely considered domain of adult ECG interpretation, whereas ECG-CPC excels in most other categories where many of the other FMs fail to reach the performance of the supervised baseline. Overall, selected ECG FMs seem promising, outperforming strong supervised baselines during finetuning and/or serving as a frozen feature extractor, achieving performance on par with supervised models during frozen evaluation. ECG-CPC, proposed in this work, is a small-scale model based on a structured state-space model backbone. It was trained on limited computational resources (a single NVIDIA L40 GPU trained for three weeks). The fact that it turned out among the strongest FMs in the benchmark suggests significant opportunities for further improving ECG foundation models. Code and model weights are provided in the supplementary material and will be made publicly available. LLMs were exclusively used for language improvement. Our framework source code and model weights for the ECG-CPC model are publicly available at https://github.com/AI4HealthUOL/ecg-fm-benchmarking.
7 Acknowledgments

### v2

@@ 4 Results
Full finetuning, frozen evaluation, and linear evaluation results are given in Tables 3, 4, and 5 (Supplementary Material). Tables indicate statistically significant differences compared to the respective best method in this task. Ranked lists for all tasks and models are provided in Table 6. We further summarize this by reporting median ranks across all tasks of a given category in Figure 2. Appendix A.4 presents comparative predictions for the 10 best-performing conditions per task, illustrating both prediction quality and label diversity.
@@ 4.1 Supervised baselines
Supervised baseline vs. literature Our supervised baseline performs on par with or surpasses the literature results. The S4-based model achieves AUROCs of 0.941 (vs. 0.9417 (Mehari and Strodthoff, 2023)) on PTB-XL, 0.908 (vs. 0.843 (Strodthoff et al., 2024)) for cardiac discharge, 0.849 (vs. 0.764 (Strodthoff et al., 2024)) for non-cardiac discharge, 0.863 (vs. 0.752 (Alcaraz et al., 2025)) for clinical deterioration, 0.747 (vs. 0.746 (Alcaraz et al., 2025)) for ICU admission, and 0.874 (vs. 0.816 (Alcaraz et al., 2025)) for mortality. Improvements on MIMIC-based tasks (except for ICU) are attributed to the multi-task training objective.
@@
Impact of model architecture We also compare the supervised S4 model with the convolutional Net1D backbone used in ECGFounder, excluding supervised transformers due to their poor performance. Table 3 reports numerical results, and Table 6 provides statistically significant rankings: S4 ranks first across tasks, while Net1D is typically four or more positions lower. Although not our primary focus, this backbone comparison reinforces prior findings that CNNs are suboptimal for physiological time series (Strodthoff et al., 2024).
4.2 Finetuning
@@
Adult ECG interpretation: Across 11 tasks on 9 datasets, ECGFounder, ECG-JEPA, and ECG-CPC are the top performers, often statistically surpassing the S4 baseline. ECG-FM ranks fourth overall, occasionally matching the leaders but underperforming on Georgia, Chapman, and PTB-XL, while MERL, ST-MEM, HuBERT-ECG, and ECGFM-KED generally fail to outperform S4. Pediatric ECG interpretation: ECG-JEPA leads, followed by ECGFounder, ST-MEM, MERL, ECG-CPC, and S4, despite the absence of pediatric pretraining data. Cardiac structure & function: ECG-CPC ranks first for echocardiography predictions, followed by ECGFounder, ECG-JEPA, ST-MEM, MERL, and S4. Cardiac and non-cardiac outcomes: ECG-CPC dominates, matching ECG-FM on non-cardiac and S4 on cardiac outcomes. ECGFounder performs relatively poorly, likely due to limited overlap between pretraining and target diagnostic labels. Acute care predictions: ECG-CPC and ECG-FM perform best across three tasks, followed by ECGFounder, ECG-JEPA, and MERL, but none significantly exceed S4. Patient characteristics: ECG-CPC ranks first in 5 of 6 tasks, surpassing S4 in 3; MERL and ECG-FM generally match or slightly trail S4, whereas ECGFounder and ECG-JEPA typically fall below it.
@@ 4.3 Frozen and linear evaluation
Figure 2: Radar plots summarizing model performance ranks (lower rank indicates statistically significantly better performance) for the eight FMs across the 7 investigated tasks. Our investigated ranking criteria accounts for confidence interval overlaps within each of the tasks and datasets. The plot is based on data from Table 7.
@@
4.4 Label efficiency
Figure 3: Scaling with dataset size on EchoNext across the best-performing FMs and supervised baseline.
Label efficiency Scaling curves for three FMs are shown in Figure 3 with scaling parameters listed in Table 30. All models stay below the supervised baseline (S4) throughout the entire range of considered training set sizes. We use the parametric form of the fits to work out a label efficiency ratio $r=N^{*}/N$, i.e., the fraction of samples $N^{*}$ required for the pretrained model to reach the same performance as the supervised baseline for given $N$. For $N$ in the range of 250 to 1000, this yields label efficiency ratios between 0.30-0.62 for ECGFounder, 0.11-0.42 for ECG-JEPA and 0.21-0.40 for ECG-CPC, see Table 31 for details. This positions ECG-JEPA as the most label-efficient model, in particular in the very low sample size regime, closely followed by ECG-CPC. This establishes the label efficiency as a relevant benchmark parameter for FMs. Furthermore, the results show that ECG FMs fulfill the promise of an improved label efficiency not only in comparison to the respective model architecture trained from scratch but also against a strong supervised baseline, improving label efficiency by up to a factor of 9. Beyond absolute label efficiency, our scaling analysis distinguishes between data-efficiency slope and performance ceiling. ECG-JEPA learns quickly with few samples (steep slope) but plateaus at a lower accuracy than ECG-CPC, which improves more slowly yet reaches a higher ceiling. Thus, FMs differ in both how fast and how far they improve as data increases, suggesting that model choice should weigh slope versus ceiling depending on whether rapid low-label learning or maximal final performance is the priority. For the case of cardiac structure analysis, this establishes ECG-JEPA as superior model in the small sample regime ($10^{3}$ training samples) and ECG-CPC as superior model in the large sample regime, thereby providing an actionable recommendation for practitioners.
4.5 Representation Similarity Analysis
We analyze intra-model CKA patterns for FMs in their original, unfinetuned form (Figure 4). ECG-CPC shows a progression from redundant early CNN layers to distinct S4 layers, with the final S4 block most specialized. ECGFounder’s convolutional stem is dissimilar to higher layers, while mid-network layers (S0-S4) are highly similar, with specialization emerging only in the final layers. In ECG-JEPA, the embedding layer is distinct, but intermediate transformer blocks (Blk1-Blk10) are nearly identical, with only the final block specialized.
(a) ECG-CPC
(b) ECGFounder
(c) ECG-JEPA
(d) ST-MEM
Figure 4: Intra-model layer-wise representation similarity analysis. CKA heatmaps comparing all internal layers within each of the best performing FM on PTB-XL (all) dataset. Higher values (yellow) indicate similar representations between layers. CKA computed using Gaussian RBF kernel ($\sigma=1.0$) on 2,500 samples per model.
ST-MEM exhibits a similar pattern, with a mostly homogeneous mid-section and specialized final layers, while the first transformer block stands out as dissimilar to neighboring blocks indicating a possible architectural bottleneck. Overall, CKA highlights that ECG-CPC exhibits the clearest and most structured evolution of representations, while other architectures show mid-layer redundancy or homogeneous transformer blocks. For a more detailed discussion including inter-model representation comparisons across depths, see Section A.9 in the supplementary material.
4.6 Model efficiency
For practical deployment, we benchmark the investigated models across GFLOP counts, GPU memory usage, and inference efficiency. Although SSM-based models are not as computationally efficient as CNNs in terms of inference efficiency, SMM (built on CPC) remains several times more efficient than transformer-based approaches across FLOPs, memory, and inference speed. More importantly, the efficiency gap between CNNs and SSMs becomes less critical when considering model quality: SMM delivers markedly stronger predictive performance than both CNNs and transformers, offering a balanced trade-off between computational cost and superior representation quality-making it an attractive choice for practical, real-world ECG deployment. The full results are given in Table 32 in the supplementary material.
@@ 5 Discussion
Details matter Reliable downstream performance depends on careful tuning. Layer-dependent learning rates consistently improve results, and some models (e.g., HuBERT-ECG, ECG-FM) fail to train at all without them. As detailed in Table 33, this benefit is strongly architecture-dependent, with transformer and SSM-based models showing marked improvements with layer-dependent learning rates. Among CNN-based models, ECGFounder remains invariant, ECGFM-KED degrades, while MERL shows modest improvements. Adjustable input size (where possible) also matters: using 2.5-second crops with test-time averaging outperforms the full 10 seconds inputs.
@@
Comparable assessment ECGFounder, ECG-JEPA, and ECG-CPC are strong on adult ECG interpretation, sometimes surpassing the supervised baseline. Pediatric ECG is dominated by ECG-JEPA. Other categories are led by ECG-CPC, followed by ECG-FM, while ECGFounder and ECG-JEPA are weaker here. Several models fall clearly below the supervised baseline, showing that self-supervised pretraining does not necessarily yield effective downstream performance. No single model consistently excels across all tasks and evaluation modes, though ECG-CPC comes closest to this goal with finetuning. This underscores that FM selection for downstream tasks should be informed by benchmarking results and similarity of the downstream task to the task categories considered here.
@@
Pretraining strategies This benchmark cannot definitively determine the optimal pretraining method, as downstream performance reflects the interaction of pretraining strategy, dataset, and model architecture (Table 1). Superior supervised performance may favor SSMs over CNNs, while transformers align naturally with masking used in self-supervised methods. Larger datasets can help (ECGFounder, ECG-CPC) but are not sufficient (e.g., HuBERT-ECG), and large-scale models do not automatically generalize. ECG-CPC’s strong performance outside ECG interpretation suggests self-supervised pretraining can be advantageous, though its unidirectional backbone limits supervised performance relative to bidirectional models. These results underscore the need for like-for-like pretraining comparisons on a common dataset (e.g., HEEDB) with standardized architectures. Fully disentangling dataset, architecture, and training strategy remains future work; our benchmark reveals practical differences, but controlled unified-dataset ablations are needed to explain them.
Model architecture In our benchmark, we provide the first systematic evidence that compact SSMs outperform much larger Transformers and comparable CNNs on time-series tasks despite having far fewer parameters, raising the theoretical question: what properties of SSM representations enable such efficient learning at minimal capacity? These models carry strong inductive biases, stable long-range memory, smooth spectral filtering, and globally parameterized convolutions, that align exceptionally well with the structure of ECG signals, enabling efficient learning even at small scale. This empirical insight, provides grounding that can catalyze new theoretical work.
S4 among SSMs Although S4 is an earlier state-space model, it remains a strong and widely used SSM baseline for continuous clinical time series due to its stable memory, spectral filtering, and computational efficiency, all of which align well with physiological signal structure. Prior ECG/EEG studies (Mehari and Strodthoff, 2023; Wang and Strodthoff, 2025) and recent SSM surveys (Somvanshi et al., 2025; Patro and Agneeswaran, 2025), consistent with our internal experiments shown in Appendix A.8, show that newer variants such as Mamba often excel in discrete domains but do not consistently outperform S4 on continuous medical signals, supporting its inclusion as a competitive and representative SSM baseline.
Model complexity The evaluated models differ markedly in size, computational efficiency, and predictive performance. Parameter counts increase from SSMs (smallest) to CNNs to transformers (largest). In terms of practical efficiency, considering GFLOPs, GPU memory, and inference latency, CNNs are fastest, followed by SSMs, with transformers being the least efficient. Importantly, SSM-based models, despite moderate computational cost, achieve the highest predictive performance, outperforming both transformers and CNNs, which show comparable accuracy.
@@
Model insights Frozen and linear evaluation reveal the nature of learned representations and the implicit knowledge captured by different FMs. Insights from probing, as used here, remain coarse, but methods from explainable AI such as analyzing representation structures and their alignment across layers and models (Vielhaben et al., 2025) could provide more detailed insights into the knowledge acquired by these models.
@@
Limitations This work has several limitations. First, the proposed tasks include only in-distribution tests; out-of-distribution evaluation would require additional datasets with compatible labels. While diagnostic label mismatches make this challenging, it is feasible for demographic data. Second, multi-task models generally improve classification performance, but for regression we applied z-normalization to mitigate issues. Still, the multi-task model used for computational efficiency may underperform single-task models on certain tasks. Third, the consistent benefit of layer-dependent learning rates during finetuning indicates potential for improved finetuning strategies. Fourth, since models are pretrained on different datasets, isolating which factors drive performance differences is difficult. Retraining on a unified dataset would enable cleaner comparisons but is computationally prohibitive. Following standard practice in vision and NLP, we prioritize practical utility by evaluating models as released; controlled unified-dataset comparisons remain important future work.
@@ 6 Conclusion
In this work, we present a comprehensive benchmark of ECG foundation models (FMs) across seven task categories. Performance varied across domains: ECGFounder, ECG-JEPA, and ECG-CPC excelled in adult ECG interpretation, while ECG-CPC outperformed others in categories where many FMs failed to match the supervised baseline. Overall, selected ECG FMs show promise, outperforming strong supervised baselines during finetuning and achieving comparable performance when used as frozen feature extractors. ECG-CPC, introduced here, is a small-scale model based on a structured state-space backbone, trained on a single NVIDIA L40 GPU for three weeks. Its strong benchmark performance highlights opportunities for further improving ECG FMs. Code and model weights are provided in the supplementary material and will be publicly released. LLMs were used only for language refinement. Our ECG-CPC framework and weights are available at https://github.com/AI4HealthUOL/ecg-fm-benchmarking.
Acknowledgments